OpenStudy (anonymous):

Prove: (can only use left side to solve right side) (sinx/1-cosx)+(1+cosx/sinx)=2(1+cosx)/sinx

OpenStudy (anonymous):

take the LCM sinx(1-cosx) and hopefully cancel values will be canceled out :)

OpenStudy (campbell_st):

well use a common denominator on the left side fractions sin(x)(1 - cos(x)) so you have \[\frac{\sin(x) \times \sin(x)}{\sin(x)(1 - \cos(x))} + \frac{(1 - \cos(x))(1 + \cos(x)}{\sin(x)(1 - \cos(x))}\] simplify the fractions and write the numeratators over the common denominator \[\frac{\sin^2(x) + 1 - \cos^2(x)}{\sin(x)(1 - \cos(x)}\]

OpenStudy (campbell_st):

so use the trig identity sin^2 = 1 - cos^2 so you have \[\frac{\sin^2(x) + \sin^2(x)}{\sin(x)(1 - \cos(x))}\] all you need to do is remove the common factor for the answer.

OpenStudy (campbell_st):

and you may need to check the right side is correctly written

OpenStudy (anonymous):

yeah its correct you get the equation (2sin^2x)/sinx(1-cosx) cancel the sinx and u get the right hand side

OpenStudy (campbell_st):

lol... my mistake with the numerator correct substitution, just wrong way around use sin^2 = 1 - cos^2 so its \[\frac{2(1 - \cos^2(x)}{\sin(x)(1 - \cos(x))}\] factor the numerator, its the difference of 2 squares... then cancel the common factor

OpenStudy (anonymous):

Thank you